OpenStudy (anonymous):

Please help and show all steps Find the equation in standard form which passes through (-1,4) and (0,9)

OpenStudy (anonymous):

slope = (y2-y1)/(x2-x1)= 5/1 =5 standard form: y=mx+b pick a point (0,9)===> 9=b y=5x+9

OpenStudy (anonymous):

m=(y1-y)/(x1-x) slope m=(9-4)/(0-(-1)) m=5/1=5 now you use ANY point to build your equation

OpenStudy (anonymous):

y-9=5(x-0) y=5x+9 standard form would look like: 5x-y+9=0